# On annihilator multiplication modules

### v1

@@ 2. Basic properties of annihilator multiplication modules
The if part follows from Corollary 2 (ii). Let $E_{i}$ be an annihilator multiplication module for each $i\in\Delta$. Choose $e\in E$. Then we can write $e=e_{i_{1}}+e_{i_{2}}+\cdots+e_{i_{n}}$ for some $i_{1},i_{2},\ldots,i_{n}\in\Delta$. Let $a\in ann(e)$. Then we conclude that $a\left(e_{i_{1}}+e_{i_{2}}+\cdots+e_{i_{n}}\right)=0$ which implies that $ae_{i_{1}}=-ae_{i_{2}}-ae_{i_{3}}-\cdots-ae_{i_{n}}\in E_{i_{1}}\cap{\textstyle\sum\limits_{j=2}^{n}}E_{i_{j}}=(0)$. Thus we conclude that $a\in ann(e_{i_{1}})$. Similarly, we have $a\in{\textstyle\bigcap\limits_{j=1}^{n}}ann(e_{i_{j}})$. Then clearly we obtain $ann(e)={\textstyle\bigcap\limits_{j=1}^{n}}ann(e_{i_{j}})$. Since $E_{i_{j}}$ is an annihilator multiplication module for each $1\leq j\leq n$, there exist finitely generated ideals $I_{j}$ such that $ann(e_{i_{j}})=ann(I_{j}E_{j})$. On the other hand, we have $ann(I_{j}E_{j})=ann(I_{j}E)$ since $ann(E_{i})=ann(E_{j})$ for every $i\neq j$ and $E$ is internal direct sum of the family $\{E_{i}\}_{i\in\Delta}$. This implies that
@@
Let $p$ be a prime number. Since $\mathbb{Q}$ is a torsion free $\mathbb{Z}$-module and $\mathbb{Z}_{p}$ is a multiplication $\mathbb{Z}$-module, so they are annihilator multiplication $\mathbb{Z}$-modules. Also, it is clear that $ann\left(\mathbb{Q})=(0\right)\neq ann(\mathbb{Z}_{p})=p\mathbb{Z}.\$However, $\mathbb{Z}$-module $E=\mathbb{Q}\oplus\mathbb{Z}_{p}$ is not an annihilator multiplication module since $ann(\overline{1})=p\mathbb{Z}\neq ann(IE)=ann(I)=(0)$ for any nonzero ideal $I$ of $\mathbb{Z}$.
Let $E$ be an $A$-module and $X$ an indeterminate over $A$. The notation $E[X]$ refers to the polynomial module over the polynomial ring $A[X]$. An $A$-module $E$ is defined as an Armendariz module if, for every $e(X)=e_{0}+e_{1}X+e_{2}X^{2}+e_{n}X^{n}$ in $E[X]$ and $f(X)=a_{0}+a_{1}X+a_{2}X^{2}+a_{k}X^{k}$ in $A[X]$ such that $f(X)e(X)=0$, the condition holds that $1\leq i\leq n$ and $1\leq j\leq k$ imply $a_{i}e_{j}=0$ [10]. Prior to examining the stability of the annihilator multiplication property in polynomial modules, the following lemma is required.
@@
Let $M$ be an Armendariz module. Then,
@@
According to [11], a proper submodule $V$ of $E$ is defined as a classical prime submodule if, whenever $abe\in V$ for some $a,b\in A$ and $e\in E$, it follows that $ae\in V$ or $be\in V$. Furthermore, $V$ is termed a classical 1-absorbing prime if, for non-units $a,b,c\in A$ and $e\in E$, the condition $abce\in V$ implies that $abe\in V$ or $ce\in V$ [36]. Every classical prime is also a classical 1-absorbing prime; however, the converse does not generally hold. The following example illustrates this distinction.
@@
Consider the $k[[X]]$-module $E=k[[X]]/(X^{2})$ where $k$ is a field and $X$ is an indeterminate over $k$. Let $V$ be the zero submodule of $E$. Then $V$ is not a classical prime submodule, since $X^{2}\overline{1}=\overline{0}\in V$ and $X\overline{1}=\overline{X}\notin V$. Let $abc\overline{e}=\overline{0}$ for some nonunits $a,b,c\in\ k[[X]]$ and $\overline{e}\in M$. Since $k[[X]]$ is a local ring with a unique maximal ideal $(X)$, we have $a=Xf$, $b=Xg$ and $c=Xh$ for some $f,g,h\in k[[X]]$. This gives $ab\overline{e}=X^{2}fg\overline{e}=\overline{0}\in V$, and so $V$ is a classical 1-absorbing prime submodule of $E$.
Recall from [34] that a proper ideal $I$ of AA\ is said to be a 1-absorbing prime ideal if whenever $abc\in I$ for some nonunits $a,b,c\in A$ then $ab\in I$ or $c\in I$.
Proposition 7.
@@
Suppose that $E$ is an annihilator multiplication module and $V$ is a submodule of $E$ such that $ann(V)\neq ann(E)$. Then there exists a proper ideal $I$ of $A$ such that $ann(V)=ann(IE)$. Let $ab\in ann(V)$ for some $a,b\in A$. Then we have $abV=aIbE=0$ which implies that $aIb\subseteq ann(E)$. If $a$ or $b$ is unit, then we are done. So assume that $a,b$ are nonunits. Since $ann(E)$ is a 1-absorbing prime ideal, we conclude that $aI\subseteq ann(E)$ or $b\in ann(E)$. Then we have $a\in ann(IE)=ann(V)$ or $b\in ann(E)\subseteq ann(V)$. Thus, $ann(V)$ is a prime ideal of $A$, so is 1-absorbing prime. Let $V,K$ be two submodules of $E$ such that $ann(V)\neq ann(E)$ and $ann(K)\neq ann(E)$. Then by above, either we have $ann(V+K)=ann(E)$ or $ann(V+K)$ is a prime ideal of $A$. Now we have two cases. Case 1: Let $ann(V+K)=ann(E)$. Since $ann(V+K)=ann(V)\cap ann(K)$, we have $ann(V)\cap ann(K)=ann(E)$. By assumption, we get $ann(K)$ and $ann(V)$ are prime ideals of $A$ which implies that $\sqrt{ann(E)}=\sqrt{ann(V)}\cap\sqrt{ann(K)}=ann(V)\cap ann(K)=ann(E)$. Since $ann(E)$ is a 1-absorbing prime ideal and semiprime, so is prime ideal. Then we have either $ann(V)\subseteq ann(K)$ or $ann(K)\subseteq ann(V)$. Case 2: Let $ann(V+K)\neq ann(E)$. In this case, $ann(V+K)=ann(V)\cap ann(K)$ is a prime ideal which implies that either $ann(V)\subseteq ann(K)$ or $ann(K)\subseteq ann(V)$. Hence, $\left\{ann(V):\ V\text{is a submodule of }E\right\}$ is totally ordered by inclusion. ∎
Proposition 8.
Proof.
The if part follows from [36, Proposition 1]. For the only if part, assume that the zero submodule is a classical 1-absorbing prime submodule. Let $abe=0$ for some $a,b\in A$ and $e\in E$. We may assume that $a,b$ are nonunits of $A$. Since $E$ is an annihilator multiplication module, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. Since $ann(e)\neq ann(E)$, $I$ is a proper ideal of $A$. This gives $abIE=0$. If $be=0$, then we are done. So, assume that $be\neq 0$. Now, we will show that $aIE=0$. Choose an arbitrary $z\in E$ and assume that $aIz\neq 0$. Since $abIz=0$, by [36, Theorem 2], we have $bz=0$. On the other hand, note that $abI(z+e)=0$. Since $aIz\neq 0$, again by [36, Theorem 2], we have $b(z+e)=0$ which implies that $be=0$ which is a contradiction. Thus, we conclude that $aIE=0$, that is, $a\in ann(IE)=ann(e)$. Then we conclude that $ae=0$ completes the proof. ∎
Remark 1.
In the previous proposition, condition $ann(e)\neq ann(E)$ is necessary. For example, consider the $k[[X]]$-module $E=k[[X]]/(X^{2})$ where $k$ is a field and $X$ is an indeterminate over $k$. Then $E$ is clearly a multiplication module, so by Example 1 it is an annihilator multiplication module. However, note that $ann(\overline{1})=(X^{2})=ann(E)$. Also, by previous example, $V=(\overline{0})$ is a classical 1-absorbing prime submodule which is not a classical prime submodule.
@@ 3. Characterizations of some special rings/modules
This section examines the characterizations of significant classes of rings and modules using the framework of annihilator multiplication modules.
@@
Let EE\ be an $A$-module. Then EE\ is a torsion free module if and only if EE\ is an annihilator multiplication module and faithful module over the domain $A.$
@@
$\left(\Longleftarrow\right):$ Let EE\ be a comultiplication and annihilator multiplication module. Then by Proposition 1 $(ii)$, for every submodule VV\ of $E$ there exists an ideal II\ of $A$ such that ann(V)=ann(IE).ann(V)=ann(IE).\ As EE\ is a comultiplication module, we have $V=ann_{E}(ann(V))=ann_{E}(ann(IE))=IE$ which completes the proof. ∎
According to [7], an $A$-module $E$ is defined as a prime module if the zero submodule of $E$ is prime, or equivalently, $ann(V)=ann(E)$ for every nonzero submodule $V$ of $E$. Additionally, a submodule $V$ of $E$ is defined to be a second submodule if, for each $a\in A$, either $aV=(0)$ or $aV=V$. In particular, $E$ is called a second module if it is a second submodule of itself [33].
An $A$-module $E$ is defined as injective if, whenever $f:L\rightarrow K$ is an injective homomorphism and $g:L\rightarrow E$ is an arbitrary homomorphism, there exists a homomorphism $h:K\rightarrow E$ such that $fh=g$ [3]. Furthermore, [33, Theorem 2.3] demonstrates that a prime module $E$ is a second module if and only if it is an injeinjective module over the ring $A/ann(E)$.
@@
Let EE\ be a second module and annihilator multiplication module. Then EE\ is a prime module and thus every submodule VV\ of EE\ is an annihilator multiplication module. In this case, VV\ is an injective $A/ann(E)$-module. In particular, if $E$ is a faithful module, then $E$ is an injective $A$-module. Furthermore, nonzero pure submodules and second submodules of EE\ coincide.
@@
Then we have ann(e)=ann(E).ann(e)=ann(E).\ Then it is clear that for every nonzero submodule VV\ of E,E,\ we have ann(V)=ann(E).ann(V)=ann(E).\ Thus, EE\ is a prime module. Furthermore, by Corollary 2 $(ii)$, VV\ is an annihilator multiplication module every nonzero submodule VV\ of E.E.\ Since $E$ is a prime and second module, by [33, Theorem 2.3], EE\ is an injective $A/ann(E)$-module. For the rest, assume that VV\ is a second submodule of E.E.\ Let JJ\ be an ideal of A.A.\ Since VV\ is a second submodule, either JV=VJV=V\ or JV=(0).JV=(0).\ If JV=(0),JV=(0),\ then $J\subseteq ann(V)=ann(E)\$which implies that $JE\cap V=(0)=JV.\$Now, assume that JV=V.JV=V.\ Then we have $JE\neq(0)\$since $ann(V)=ann(E)$. As EE\ is a second module, we have JE=EJE=E\ which implies that $JE\cap V=V=JV.\$Thus, VV\ is a pure submodule. For the converse, assume that VV\ is a nonzero pure submodule of EE\ and $a\in A.\$Then we have $aV=aE\cap V.\$Without loss of generality, we may assume that $aV\neq 0.\$As $ann(V)=ann(E)$ and EE\ is second,,\ we have $aE\neq(0)\$and so aE=E.aE=E.\ This gives $aE\cap V=V=aV,\$that is, VV\ is a second submodule of E.E.\
It is well known that a ring $A$ is von Neumann regular if and only if $A$ is reduced (i.e, a ring without a nonzero nilpotent) and every prime ideal is maximal. Now, we characterize principal ideal von Neumann regular rings in terms of annihilator multiplication modules.
@@
(iii) AA\ is a direct product of finitely many fields.

### v2

@@ 2. Basic properties of annihilator multiplication modules
The if part follows from Corollary 2 (ii). For the only if part, let $E_{i}$ be an annihilator multiplication module for each $i\in\Delta$. Choose $e\in E$. Then we can write $e=e_{i_{1}}+e_{i_{2}}+\cdots+e_{i_{n}}$ for some $i_{1},i_{2},\ldots,i_{n}\in\Delta$. Let $a\in ann(e)$. Then we conclude that $a\left(e_{i_{1}}+e_{i_{2}}+\cdots+e_{i_{n}}\right)=0$ which implies that $ae_{i_{1}}=-ae_{i_{2}}-ae_{i_{3}}-\cdots-ae_{i_{n}}\in E_{i_{1}}\cap{\textstyle\sum\limits_{j=2}^{n}}E_{i_{j}}=(0)$. Thus we conclude that $a\in ann(e_{i_{1}})$. Similarly, we have $a\in{\textstyle\bigcap\limits_{j=1}^{n}}ann(e_{i_{j}})$. Then clearly we obtain $ann(e)={\textstyle\bigcap\limits_{j=1}^{n}}ann(e_{i_{j}})$. Since $E_{i_{j}}$ is an annihilator multiplication module for each $1\leq j\leq n$, there exist finitely generated ideals $I_{j}$ such that $ann(e_{i_{j}})=ann(I_{j}E_{j})$. On the other hand, we have $ann(I_{j}E_{j})=ann(I_{j}E)$ since $ann(E_{i})=ann(E_{j})$ for every $i\neq j$ and $E$ is internal direct sum of the family $\{E_{i}\}_{i\in\Delta}$. This implies that
@@
Let $p$ be a prime number. Since $\mathbb{Q}$ is a torsion-free $\mathbb{Z}$-module and $\mathbb{Z}_{p}$ is a multiplication $\mathbb{Z}$-module, so they are annihilator multiplication $\mathbb{Z}$-modules. Also, it is clear that $ann\left(\mathbb{Q})=(0\right)\neq ann(\mathbb{Z}_{p})=p\mathbb{Z}.\$However, $\mathbb{Z}$-module $E=\mathbb{Q}\oplus\mathbb{Z}_{p}$ is not an annihilator multiplication module since $ann(\overline{1})=p\mathbb{Z}\neq ann(IE)=ann(I)=(0)$ for any nonzero ideal $I$ of $\mathbb{Z}$.
Let $E$ be an $A$-module and $X$ an indeterminate over $A$. The notation $E[X]$ refers to the polynomial module over the polynomial ring $A[X]$. An $A$-module $E$ is defined as an Armendariz module if, for every $e(X)=e_{0}+e_{1}X+e_{2}X^{2}+\cdots+e_{n}X^{n}$ in $E[X]$ and $f(X)=a_{0}+a_{1}X+a_{2}X^{2}+\cdots+a_{k}X^{k}$ in $A[X]$ such that $f(X)e(X)=0$, then $a_{i}e_{j}=0$ for every $1\leq i\leq k$ and $1\leq j\leq n$ [14]. Prior to examining the stability of the annihilator multiplication property in polynomial modules, the following lemma is required.
@@
Let $E$ be an Armendariz module. Then,
@@
Let $E$ be an $A$-module. The trivial extension $A\propto E=A\oplus E$ of $A$-module $E$ is a commutative ring with a nonzero identity $(1,0)$ under componentwise addition and following multiplication: $(a,e)(b,v)=(ab,av+be)$ for each $(a,e),(b,v)\in A\propto E$ (See, [6] and [23]). For an ideal $I$ of $A$ and a submodule $V$ of $E$, $I\propto V$ is an ideal of $A\propto E$ if and only if $IE\subseteq V$ [6, Theorem 3.1]. Now, we are ready to determine certain annihilator multiplication/multiplication ideals in trivial ring extension $A\propto E$.
Theorem 1.
Let $E$ be a finitely generated $A$-module and $V$ a submodule of $E$. The following statements are equivalent.
  • (i)
    $V$ is an annihilator multiplication module.
  • (ii)
    $0\propto V$ is an annihilator multiplication ideal of $A\propto E$.
Proof.
$\left(i\right)\Leftarrow\left(ii\right):$ Let $V$ be an annihilator multiplication module and choose $(0,v)\in\ 0\propto V$. Since $V$ is an annihilator multiplication module, there exists a finitely generated ideal $I$ of $A$ such that $ann(v)=ann(IV)$. This gives $ann(0,v)=ann(v)\propto E=ann(IV)\propto E=ann((I\propto E)(0\propto V))$. As $I$ and $E$ are finitely generated, it follows that $I\propto E$ is a finitely generated ideal in $A\propto E$. Thus, $0\propto V$ is an annihilator multiplication ideal in $A\propto E$.
$\left(ii\right)\Leftarrow\left(i\right):$ Let $0\propto V$ be an annihilator multiplication ideal of $A\propto E$. Choose $v\in V$, and by the assumption, note that $ann(0,v)=ann(\mathfrak{J}(0\propto V))$ for some finitely generated ideal $\mathfrak{J}$ of $A\propto E$. As $E$ is finitely generated, it follows that $\mathfrak{J}+(0\propto E)$ is also a finitely generated ideal in $A\propto E$. Since $(\mathfrak{J}+0\propto E)(0\propto V)=\mathfrak{J}(0\propto V)+(0\propto E)(0\propto V)=\mathfrak{J}(0\propto V)$, we may assume that $\mathfrak{J}$ contains $0\propto E$. In this case, by [6, Theorem 3.1], $\mathfrak{J}=I\propto E$ and $I$ is a finitely generated ideal in $A$. Thus we have $\mathfrak{J}(0\propto V)=0\propto IV$, and this gives $ann(0,v)=ann(v)\propto E=ann(\mathfrak{J}(0\propto V))=ann(0\propto IV)=ann(IV)\propto E$. Therefore, we conclude that $ann(v)=ann(IE)$ which completes the proof. ∎
Recall from [38] that an $A$-module $E$ is said to be a simple module if $Ae=E$ for every $0\neq e\in E$. Now, we are ready to characterize simple modules in terms of certain multiplication ideals of the trivial extension $A\propto E$.
Theorem 2.
Let $A$ be a local ring with unique maximal ideal $\mathfrak{m}$ and $E$ be an $A$-module such that $\mathfrak{m}E=0$. Suppose that $V$ is a submodule of $E$. Then $0\propto V$ is a multiplication ideal of $A$ if and only if $V$ is a simple module.
Proof.
$(\Rightarrow):$ Suppose that $0\propto V$ is a multiplication ideal of $A\propto E$. Choose $0\neq v\in V$. Since $0\propto V$ is a multiplication ideal and $(0,v)\in 0\propto V$. By [19, Lemma 3.5], there exists $(a_{1},v_{1}),(a_{2},v_{2}),\ldots,(a_{n},v_{n})\in A\propto E$ such that $A\propto E(0,v)=0\propto Av=\sum_{i=1}^{n}(a_{i},v_{i})0\propto V$. If $a_{i}$’s are nonunits of $A$, then $a_{i}\in\mathfrak{m}$. Then by the assumption, $\sum_{i=1}^{n}(a_{i},v_{i})0\propto V=0\propto 0=0\propto Av$ which implies that $Av=0$, a contradiction. Then there exists $i=1,2,\ldots,n$ such that $a_{i}$ is a unit of $A$. In this case, we have $\sum_{i=1}^{n}(a_{i},v_{i})0\propto V=0\propto V=A\propto E(0,v)=0\propto Av$ which implies that $Av=V$, that is $V$ is a simple module.
$(\Leftarrow):$ Assume that $V$ is a simple module. Choose a nonzero element $(0,v)\in 0\propto V$. Since $V$ is a simple module and $v\neq 0$, we have $Av=V$. This implies that $A\propto E(0,v)=0\propto Av=0\propto V=(A\propto E)(0\propto V)$, that is $0\propto V$ is a multiplication ideal of $A\propto E$. ∎
Theorem 3.
Let $A$ be a local ring with unique maximal ideal $\mathfrak{m}$ and $E$ be an $A$-module such that $\mathfrak{m}E=0$. For a proper ideal $I$ of $A$, $I\propto E$ is an annihilator multiplication ideal of $A\propto E$ if and only if $I$ is an annihilator multiplication ideal of $A$.
Proof.
$(\Leftarrow):$ Suppose that $I$ is an annihilator multiplication ideal of $A$ and choose $(a,v)\in I\propto E$. Then $a\in I\subseteq\mathfrak{m}$ which implies that $ann(a,v)=ann(a)\propto E$. Since $I$ is an annihilator multiplication ideal and $a\in I$, there exists a finitely generated ideal $J=\sum_{i=1}^{k}Ax_{i}$ of $A$ such that $ann(a)=ann(JI)=\bigcap_{i=1}^{k}ann(x_{i}I)$. Now, we have two cases. Case 1: Assume that $J=A$. In this case, we obtain $ann(a)=ann(I)$ which implies that $ann(a,v)=ann(a)\propto E=ann(I)\propto E=ann((1,0)I\propto E)$. Thus, $I\propto E$ is an annihilator multiplication ideal of $A\propto E$. Case 2: Suppose that $J$ is a proper ideal of $A$. In this case, $x_{i}$ is a nonunit and hence an element of $\mathfrak{m}$ for each $i=1,2,\ldots,k$. Since $\mathfrak{m}E=0$, it is easy to see that $(x_{i},0)I\propto E=x_{i}I\propto 0$ which implies that $ann(x_{i}I)\propto E=ann((x_{i},0)I\propto E)$. Then we conclude that $ann(a,v)=ann(a)\propto E=(\bigcap_{i=1}^{k}ann(x_{i}I))\propto E=\bigcap_{i=1}^{k}(ann(x_{i}I)\propto E)=\bigcap_{i=1}^{k}ann((x_{i},0)I\propto E)$. Now, put $\mathfrak{J}=\sum_{i=1}^{k}(x_{i},0)A\propto E$. Then $\mathfrak{J}$ is a finitely generated ideal in $A\propto E$ and it is clear that $ann(a,v)=ann(\mathfrak{J}(I\propto E))$ which completes the proof.
$(\Rightarrow):$ Choose an element $a\in I$. Since $I\propto E$ is an annihilator multiplication ideal and $\mathfrak{m}E=0$, we have $ann(a,0_{E})=ann(a)\propto E=ann(\mathfrak{J}(I\propto E))$ for some finitely generated ideal $\mathfrak{J}$ of $A\propto E$. Now, we have two cases. Case 1: Let $\mathfrak{J}=A\propto E$. Then we have $ann(a,0_{E})=ann(a)\propto E=ann(I\propto E)=ann(I)\propto E$ which implies that $ann(a)=ann(I)$. As $A$ is principal ideal and $AI=I$, it follows that $I$ is an annihilator multiplication ideal of $A$. Case 2: Let $\mathfrak{J}$ be a proper ideal of $A\propto E$. Then for any $(b,v)\in\mathfrak{J}$, $b\in\mathfrak{m}$ because $b\in A$ is nonunit and $\mathfrak{m}$ is the unique maximal ideal of $A$. This gives that $(b,v)(0,z)=(0,bz)=(0,0)$ for any $z\in E$, and thus we obtain $\mathfrak{J}(0\propto E)=0\propto 0$. Then we conclude that $\mathfrak{J}[(I\propto E)+(0\propto E)]=\mathfrak{J}(I\propto E)$. Thus, we may assume that $\mathfrak{J}$ contains $0\propto E$. Then by [6, Theorem 3.1], we can write $\mathfrak{J}=K\propto E$ for some finitely generated proper ideal $K$ of $A$. This implies that $\mathfrak{J}(I\propto E)=(K\propto E)(I\propto E)=KI\propto 0$. Then we have $ann(a,0_{E})=ann(a)\propto E=ann(\mathfrak{J}(I\propto E))=ann(KI\propto 0)=ann(KI)\propto E$ which implies that $ann(a)=ann(KI)$ for some finitely generated ideal $K$ of $A$, that is $I$ is an annihilator multiplication ideal.
The following example provides a construction of an annihilator multiplication ring that is not a multiplication ring.
@@
Let $E$ be a vector space over a field $k$ with $dim_{k}(E)=n\geq 2$. Consider the trivial extension $R=k\propto E$. Then $R$ is an annihilator multiplication ring which is not a multiplication ring.
@@
Suppose that $R=k\propto E$, where $E$ is a vector space over the field $k$ with $dim_{k}(E)=n\geq 2$. Then by [6, Corollary 3.4], we know that every proper ideal of $R$ has the form $0\propto V$ for some subspace $V$ of $E$. Since $E$ is a vector space, its all subspaces are torsion-free, so by Example 1, $V$ is an annihilator multiplication module. Then by Theorem 1, $0\propto V$ is an annihilator multiplication ideal of $R$. Thus, $R$ is an annihilator multiplication ring. On the other hand, put $V=E$. Since $E$ is not a simple module, by Theorem 2, $0\propto V$ is not a multiplication ideal of $R$, and hence $R$ can not be a multiplication ring. ∎
@@ 3. Characterizations of some special rings/modules
This section is devoted to characterizing several significant classes of rings and modules, including torsion-free, multiplication, uniserial, injective modules, classical prime/classical 1-absorbing prime submodules, and Noetherian vn-regular rings. Furthermore, we investigate the relationship between the associated primes of the module, $Ass_{A}(E)$, and the associated primes of the underlying ring, $Ass(A)$, within the framework of annihilator multiplication modules.
@@
Let EE\ be an $A$-module. Then EE\ is a torsion-free module if and only if EE\ is an annihilator multiplication module and faithful module over the domain $A.$
@@
$\left(\Leftarrow\right):$ Let EE\ be a comultiplication and annihilator multiplication module. Then by Proposition 1 $(ii)$, for every submodule VV\ of $E$ there exists an ideal II\ of $A$ such that ann(V)=ann(IE).ann(V)=ann(IE).\ As EE\ is a comultiplication module, we have $V=ann_{E}(ann(V))=ann_{E}(ann(IE))=IE$ which completes the proof. ∎
According to [15], a proper submodule $V$ of $E$ is defined as a classical prime submodule if, whenever $abe\in V$ for some $a,b\in A$ and $e\in E$, it follows that $ae\in V$ or $be\in V$. Furthermore, $V$ is called a classical 1-absorbing prime if, for some nonunits $a,b,c\in A$ and $e\in E$, the condition $abce\in V$ implies that $abe\in V$ or $ce\in V$ [45]. One can easily see that every classical prime is also a classical 1-absorbing prime; however, the converse does not generally hold. The following example illustrates this distinction.
Example 7.
Consider the $k[[X]]$-module $E=k[[X]]/(X^{2})$ where $k$ is a field and $X$ is an indeterminate over $k$. Let $V$ be the zero submodule of $E$. Then $V$ is not a classical prime submodule, since $X^{2}\overline{1}=\overline{0}\in V$ and $X\overline{1}=\overline{X}\notin V$. Let $abc\overline{e}=\overline{0}$ for some nonunits $a,b,c\in\ k[[X]]$ and $\overline{e}\in E$. Since $k[[X]]$ is a local ring with a unique maximal ideal $(X)$, we have $a=Xf$, $b=Xg$ and $c=Xh$ for some $f,g,h\in k[[X]]$. This gives $ab\overline{e}=X^{2}fg\overline{e}=\overline{0}\in V$, and so $V$ is a classical 1-absorbing prime submodule of $E$.
@@
Proof.
The if part follows from [45, Proposition 1]. For the only if part, assume that the zero submodule is a classical 1-absorbing prime submodule. Let $abe=0$ for some $a,b\in A$ and $e\in E$. We may assume that $a,b$ are nonunits of $A$. Since $E$ is an annihilator multiplication module, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. Since $ann(e)\neq ann(E)$, $I$ is a proper ideal of $A$. This gives $abIE=0$. If $be=0$, then we are done. So, assume that $be\neq 0$. Now, we will show that $aIE=0$. Choose an arbitrary $z\in E$ and assume that $aIz\neq 0$. Since $abIz=0$, by [45, Theorem 2], we have $bz=0$. On the other hand, since $abIe=0$ and $be\neq 0$, we have $aIe=0$. As $abI(z+e)=0$ and $b(z+e)=be\neq 0$, we conclude that $aI(z+e)=aIz=0$ which is a contradiction. Thus, we conclude that $aIE=0$, that is, $a\in ann(IE)=ann(e)$. This gives that $ae=0$, as required. ∎
Remark 1.
In the previous proposition, the condition $ann(e)\neq ann(E)$ is necessary. For example, consider the $k[[X]]$-module $E=k[[X]]/(X^{2})$, where $k$ is a field and $X$ is an indeterminate over $k$. Then $E$ is clearly a multiplication module, so by Example 1, it is an annihilator multiplication module. However, note that $ann(\overline{1})=(X^{2})=ann(E)$. Also, by Example 7, $V=(\overline{0})$ is a classical 1-absorbing prime submodule which is not a classical prime submodule.
Recall from [43] that a proper ideal $I$ of AA\ is said to be a 1-absorbing prime ideal if whenever $abc\in I$ for some nonunits $a,b,c\in A$, then $ab\in I$ or $c\in I$.
Proposition 9.
Proof.
Suppose that $E$ is an annihilator multiplication module and $V$ is a submodule of $E$ such that $ann(V)\neq ann(E)$. Then there exists a proper ideal $I$ of $A$ such that $ann(V)=ann(IE)$ by Proposition 1 (ii). Let $ab\in ann(V)$ for some $a,b\in A$. Then we have $abV=aIbE=0$ which implies that $aIb\subseteq ann(E)$. If $a$ or $b$ is unit, then we are done. So assume that $a,b$ are nonunits. Since $ann(E)$ is a 1-absorbing prime ideal, we conclude that $aI\subseteq ann(E)$ or $b\in ann(E)$. Then we have $a\in ann(IE)=ann(V)$ or $b\in ann(E)\subseteq ann(V)$. Thus, $ann(V)$ is a prime ideal of $A$, so is 1-absorbing prime. Let $V,K$ be two submodules of $E$ such that $ann(V)\neq ann(E)$ and $ann(K)\neq ann(E)$. Then by above, either we have $ann(V+K)=ann(E)$ or $ann(V+K)$ is a prime ideal of $A$. Now we have two cases. Case 1: Let $ann(V+K)=ann(E)$. Since $ann(V+K)=ann(V)\cap ann(K)$, we have $ann(V)\cap ann(K)=ann(E)$. By assumption, we get $ann(K)$ and $ann(V)$ are prime ideals of $A$ which implies that $\sqrt{ann(E)}=\sqrt{ann(V)}\cap\sqrt{ann(K)}=ann(V)\cap ann(K)=ann(E)$. Since $ann(E)$ is a 1-absorbing prime ideal and semiprime, so is prime ideal. Then we have either $ann(V)\subseteq ann(K)$ or $ann(K)\subseteq ann(V)$. Case 2: Let $ann(V+K)\neq ann(E)$. In this case, $ann(V+K)=ann(V)\cap ann(K)$ is a prime ideal which implies that either $ann(V)\subseteq ann(K)$ or $ann(K)\subseteq ann(V)$. Hence, $\left\{ann(V):\ V\ \text{is a submodule of }E\right\}$ is totally ordered by inclusion. ∎
Let $E$ be an $A$-module. Recall from [21] that $E$ is said to be a uniserial module if its each two submodules are comparable, that is, $V\subseteq W$ or $W\subseteq V$ for every two submodules $V,W$ of $E$. For more details on uniserial modules, we refer [21] and [39] to the reader.
Corollary 3.
Let $E$ be an annihilator multiplication module and a comultiplication module such that $Ann(E)$ is a 1-absorbing prime ideal of $A$. Then $E$ is a uniserial module.
Proof.
Suppose that $E$ is an annihilator multiplication module and a comultiplication module. Let $V,W$ be two submodules of $E$. Then by Proposition 9, we have either $ann(V)\subseteq ann(W)$ or $ann(W)\subseteq ann(V)$. Without loss of generality, we may assume that $ann(V)\subseteq ann(W)$. Since $E$ is a comultiplication module, we have $W=ann_{E}(ann(W))\subseteq ann_{E}(ann(V))=V$, that is $E$ is a uniserial module. ∎
According to [9], an $A$-module $E$ is defined as a prime module if the zero submodule of $E$ is prime, or equivalently, $ann(V)=ann(E)$ for every nonzero submodule $V$ of $E$. Additionally, a submodule $V$ of $E$ is defined to be a second submodule if, for each $a\in A$, either $aV=(0)$ or $aV=V$. In particular, $E$ is called a second module if it is a second submodule of itself [42]. An $A$-module $E$ is defined as injective if, whenever $f:L\rightarrow K$ is an injective homomorphism and $g:L\rightarrow E$ is an arbitrary homomorphism, there exists a homomorphism $h:K\rightarrow E$ such that $fh=g$ [7]. Furthermore, [42, Theorem 2.3] demonstrates that a prime module $E$ is a second module if and only if it is an injective module over the ring $A/ann(E)$.
Proposition 10.
Let EE\ be a second module and annihilator multiplication module. Then EE\ is a prime module and thus every submodule VV\ of EE\ is an annihilator multiplication module. In this case, EE\ is an injective $A/ann(E)$-module. In particular, if $E$ is a faithful module, then $E$ is an injective $A$-module. Furthermore, nonzero pure submodules and second submodules of EE\ coincide.
@@
Then we have $ann(e)=ann(E)$. This implies that $ann(V)=ann(E)$ for every nonzero submodule $V$ of $E$. Thus, EE\ is a prime module. Furthermore, by Corollary 2 $(ii)$, VV\ is an annihilator multiplication module for every nonzero submodule VV\ of E.E.\ Since $E$ is prime and second module, by [42, Theorem 2.3], EE\ is an injective $A/ann(E)$-module. For the rest, assume that VV\ is a second submodule of E.E.\ Let JJ\ be an ideal of A.A.\ Since VV\ is a second submodule, either JV=VJV=V\ or JV=(0).JV=(0).\ If JV=(0),JV=(0),\ then $J\subseteq ann(V)=ann(E)\$which implies that $JE\cap V=(0)=JV.\$Now, assume that JV=V.JV=V.\ Then we have $JE\neq(0)\$since $ann(V)=ann(E)$. As EE\ is a second module, we have JE=EJE=E\ which implies that $JE\cap V=V=JV.\$Thus, VV\ is a pure submodule. For the converse, assume that VV\ is a nonzero pure submodule of EE\ and $a\in A.\$Then we have $aV=aE\cap V.\$Without loss of generality, we may assume that $aV\neq 0.\$As $ann(V)=ann(E)$ and EE\ is second,,\ we have $aE\neq(0)\$and so aE=E.aE=E.\ This gives $aE\cap V=V=aV,\$that is, VV\ is a second submodule of E.E.\
It is well known that a ring $A$ is von Neumann regular if and only if $A$ is reduced (i.e, a ring without a nonzero nilpotent) and every prime ideal is maximal. Now, we characterize Noetherian von Neumann regular rings in terms of annihilator multiplication modules.
@@
(iii) $A$ is a Noetherian von Neumann regular ring.
(iv) AA\ is a direct product of finitely many fields.